OpenStudy (bookworm14):

Help Please! solve the following equation: 8+2x+6=3x+12 I'm not sure how to work this one out, any help would be great THANKS!!!- Bree :)

OpenStudy (anonymous):

@Bookworm14 8 + 2x + 6 = 3x + 12 2x + 14 = 3x + 12 2x + 14 - 3x = 3x + 12 - 3x x + 14 = 12 x + 14 - 14 = 12 - 14 -x = -2 -x/-1 = -2/-1 x = 2

OpenStudy (bookworm14):

wouldn't 2x-3x be -x , instead of just x? (look in line 2-3)

OpenStudy (anonymous):

8 + 2x + 6 = 3x + 12 8 + 6 - 12 = 3x - 2x 2 = x

OpenStudy (anonymous):

@Bookworm14 got it??

OpenStudy (bookworm14):

ok I think I get it, thanks yall !!- Bree :)

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

\[x2_3x_9divx+1\]